OpenStudy (anonymous):

how to integrate log | 4x^2-1| dx ?

OpenStudy (anonymous):

start by factoring

OpenStudy (anonymous):

then break apart in to two pieces then integrate term by term

Elsa213 (elsa213):

erm.... integral 4x/(1+x^2) = I Let 1+x^2 = t ...........(1) --> 2x = dt/dx --> 2x dx = dt ...........(2) Now I = integral 2 [2x dx / (1+x^2) ] = integral 2 [dt / t ] since from (2) = 2 [integral dt / t ] = 2 (ln t) +c = 2 ln(1+x^2) + c We have integral 1/t dt = ln t +c Therefore I = 2 ln(1+x^2) +c where ln is log to the base e https://answers.yahoo.com/question/index?qid=20110416060345AAVswO7

OpenStudy (anonymous):

@Elsa213 i am pretty sure that is some other answer to some other question

Elsa213 (elsa213):

ik ik

Elsa213 (elsa213):

it was just an example

OpenStudy (anonymous):

log | (2x+1)(2x-1) | ??

OpenStudy (xapproachesinfinity):

log a.b=loga +logb then integrate each at a time

OpenStudy (anonymous):

right

OpenStudy (anonymous):

now \[\int \log(2x+1)dx+\int (2x-1)dx\]

OpenStudy (xapproachesinfinity):

you forgot log lol

OpenStudy (anonymous):

then you should be done because you should have the integral of the log memorized

OpenStudy (anonymous):

so i did \[\int \log(2x+1)dx+\int \log(2x-1)dx\]

OpenStudy (anonymous):

the answer given is : x log | 4x^2-1| - 1/2 log | (2x-1)/(2x+1) | .. will i get this answer this way ?

OpenStudy (xapproachesinfinity):

yes continue this way

OpenStudy (anonymous):

ok thanks @xapproachesinfinity and @satellite73

OpenStudy (xapproachesinfinity):

you should know what is int log|x|dx

OpenStudy (anonymous):

i have no idea when you integrate you will get something then using the properties of the log you can change it in to many different forms don't get married to the answer in the book

OpenStudy (anonymous):

because this exercise was about the problems of integral dx / (a^2_x^2) type

OpenStudy (anonymous):

i would say this integral looks nothing like that at all

OpenStudy (xapproachesinfinity):

satellite you assumed that the expression will take positive values only you need the absolute value theere

OpenStudy (xapproachesinfinity):

i would say this integral looks nothing like that at all === is this to my reply^_^

OpenStudy (anonymous):

yes the | 4x^2 -1 | was there

OpenStudy (anonymous):

i would say \[\int \log(4x^2-1)dx\] looks nothing like \[\int \frac{dx}{a^2-x^2}\]

OpenStudy (xapproachesinfinity):

yes i doesn't :)

OpenStudy (xapproachesinfinity):

it*

OpenStudy (anonymous):

or dx/x^2-a^2

OpenStudy (anonymous):

\[\int \log(x)dx=x\log(x)-x\] so \[\int \log(2x+1)dx=\frac{1}{2}\left((2x+1)\log(2x-1)-(2x+1)\right)\] by a mental u sub second part is similar

OpenStudy (xapproachesinfinity):

@satellite73 , any reason why you took off absolute values?

OpenStudy (anonymous):

damn typo \[\int \log(2x+1)dx=\frac{1}{2}\left((2x+1)\log(2x+1)-(2x+1)\right)\]

OpenStudy (anonymous):

or if you prefer \[\frac{1}{2}(2x+1)(\log(2x+1)-1)\]

OpenStudy (anonymous):

laziness is all that and the stupid \(+ c\) got omitted as well

OpenStudy (xapproachesinfinity):

but you excluded the negative values?

ganeshie8 (ganeshie8):

why wolfram is giving this mess http://www.wolframalpha.com/input/?i=%5Cint+log%28abs%284x%5E2-1%29%29

OpenStudy (xapproachesinfinity):

wolfram is crazy hehe

OpenStudy (xapproachesinfinity):

i say the absolute value cannot be taken off. as we are excluding negative interval here

OpenStudy (anonymous):

wolfram gave that mess because of the sign problem here it is neater http://www.wolframalpha.com/input/?i= \int+log%284x^2-1%29

OpenStudy (xapproachesinfinity):

\(\Large \rm\color{midnightblue}{\int log|(2x+1)(2x-1)|dx=\int log|2x+1|dx+\\ \int log|2x-1|dx}\)

myininaya (myininaya):

Integration by parts should even work for the first way this expression appears

myininaya (myininaya):

integration by parts will be needed polynomial division partial fractions look like they are needed as well

OpenStudy (xapproachesinfinity):

i agree. i looks that the answer would look much simpler

OpenStudy (xapproachesinfinity):

x/(2x-1)(2x+1) partial fraction here

OpenStudy (xapproachesinfinity):

the other part is good xlog|4x^2-1|

myininaya (myininaya):

Whenever I see something that looks like \[\int\limits_{}^{}1 \cdot \ln(f(x)) dx \] one of my first instincts it to go with the integration by parts (assuming f>0) \[=x \ln(f(x))-\int\limits_{}^{}x \frac{f'(x)}{f(x)} dx\]

OpenStudy (xapproachesinfinity):

so far it looks nice to me

myininaya (myininaya):

where did you get that @xapproachesinfinity ?

OpenStudy (xapproachesinfinity):